# revtex4-1Repair the float

### v1

@@ Acknowledgements.
L. L. and B. H. acknowledge funding from the National Science Foundation under Grant No. 2145147. M. Jr. Aguirre acknowledges funding from the University of California, Riverside. This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-21-1-0263. B. K. K. acknowledges that part of this work was done under the auspices of the U.S. Department of Energy under Project No. 20240256ER of the Laboratory Directed Research and Development Program at Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy (Contract No. 89233218CNA000001).

### v2

@@ Acknowledgements.
L. L. , M. A. Jr. and B. H. acknowledge funding from the National Science Foundation under Grant No. 2145147. M. A. Jr. acknowledges funding from the University of California, Riverside. This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-21-1-0263. B. K. K. acknowledges that part of this work was done under the auspices of the U.S. Department of Energy under Project No. 20240256ER of the Laboratory Directed Research and Development Program at Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy (Contract No. 89233218CNA000001).
Data Availability
The data that support the findings of this article are openly available Liu et al. (2025).